OpenStudy (anonymous):

Determine if the equation Ax = b is consistent for all possible b1 and b2 . If the equation is not consistent for all possible b1 and b2 , give a description of the set of all b for which the equation is consistent (i.e., the condition which must be satisfied by b1 and b2 )

OpenStudy (anonymous):

1-The equation is consistent for all possible b1 and b2 . 2-The equation is consistent for all b1 and b2 satisfying b1 + b2 =0. 3-The equation is consistent for all b1 and b2 satisfying b1 - 8b2 =0. 4-The equation is consistent for all b1 and b2 satisfying 8b1 - b2 =0. 5-None of the above.

OpenStudy (anonymous):

@ganeshie8 could you please help me with this one?

OpenStudy (anonymous):

im not sure what i have to do to the matrix to work this out

OpenStudy (anonymous):

@sidsiddhartha could you please help me with this one?

OpenStudy (phi):

you can show it is full rank by showing it has 2 pivots. (if that makes any sense)

ganeshie8 (ganeshie8):

Notice that the matrix has `two independent column vectors` - can you reach all the points in a `plane` using these two column vectors ?

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

what do you mean by two independent column vectors?

ganeshie8 (ganeshie8):

they are pointing in different directions, so they're linearly independent column vectors : you cannot get one vector from the other by scaling

OpenStudy (anonymous):

ohk!

OpenStudy (anonymous):

so no you cant reach all the points? using just those two vectors

ganeshie8 (ganeshie8):

why not ? tell me one point which you think you cannot reach using linear combinations of these two vectors

OpenStudy (anonymous):

oh using linear combination of these vectors i suppose u can

ganeshie8 (ganeshie8):

yes you can reach all the points by taking linear combinations of the columns of \(A\), that means the equation \(Ax=b\) has solutions for all \(b\)

OpenStudy (haseeb96):

\[\left[\begin{matrix}-8 & 8 \\ 8 & 8\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}b1 \\ b2\end{matrix}\right)\]

ganeshie8 (ganeshie8):

a fancy word for saying the same is \(\text{consistent}\)

OpenStudy (haseeb96):

according to yur condition Ax=b

OpenStudy (anonymous):

sorry nothing is showing up

OpenStudy (haseeb96):

\[\left[\begin{matrix}-8x+8y \\ 8x+8y\end{matrix}\right] = \left(\begin{matrix}b1 \\ b2\end{matrix}\right)\]

OpenStudy (haseeb96):

but here it is showing just put x=IxI IyI

OpenStudy (haseeb96):

u just refresh yur page

OpenStudy (haseeb96):

 this is that thing which was written there .